OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = x2 + 2x – 3? Answer a.x–intercepts: (3, 0) and (–1, 0) y–intercept: (0, –3) b.x–intercepts: (–3, 0) and (1, 0) y–intercept: (0, 3) c.x–intercepts: (–3, 0) and (1, 0) y–intercept: (0, –3) d.x–intercepts: (3, 0) and (–1, 0) y–intercept: (0, 3)

OpenStudy (anonymous):

Factor to get y=(x+3)(x-1) Set y=0 and solve to get x+3=0 or x-1=0 x=-3 or x=1 These are the x-intercepts

OpenStudy (anonymous):

y-int plug in 0 for x y = -3 x-int plug in 0 for y 0 = x^2 + 2x - 3 factor 0 = (x-1)(x+3) x-int: x = 1, -3